Carnes Cosmetics Co.'s stock price is $54.66, and it recently paid a $1.25 dividend. This dividend is expected to grow by 22% for the next 3 years, then grow forever at a constant rate, g; and rs = 10%. At what constant rate is the stock expected to grow after Year 3? Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

D0 =1.25
D1 =1.25*(1+22%)
D2 =1.25*(1+22%)^2
D3=1.25*(1+22%)^3
D4 =1.25*(1+22%)^3*(1+g)
Terminal Value =1.25*(1+22%)^3*(1+g)/(10%-g)
Price of Stock =54.66 =1.25*(1+22%)/1.10+1.25*(1+22%)^2/1.10^2+1.25*(1+22%)^3/1.10^3+1.25*(1+22%)^3*(1+g)/((10%-g)*1.10^3)
54.66 =4.6293+1.7053418(1+g)/(0.10-g)
(54.66-4.6293)/1.7053418 =(1+g)/(0.1-g)
28.3376317841582=(1+g)/(0.1-g)
28.3376317841582*0.1-28.3376317841582*g =1+g
g =6.82%